OpenStudy (anonymous):

what is the formula for solving the following word problem? In Allison's class, there are 23 more boys than girls. If there are a total of 81 students in the class, how many girls are in the class?

OpenStudy (anonymous):

Define some variables. B = number of boys G = number of girls B = G+23

OpenStudy (anonymous):

thank you ----I've tried backsolving but wanted a more algebraic approach--52 +n=81 but I get stuck from that point.

OpenStudy (anonymous):

I see this is a great help.

OpenStudy (pfenn1):

You know B=G + 23 You also know the number of girls and boys together = 81 B + G=81 substitute in for B and solve for G